# Macroscopic degeneracy of ground state in the frustrated Heisenberg diamond chain

### v1

@@ I Introduction
The low-dimensional quantum magnets on geometrically frustrated lattices have been extensively studied during last years diep ; mila . An interesting class of the frustrated quantum magnets are models in which the ground state is macroscopically degenerate. The systems with macroscopic degeneracy have attracted great interest. One particular class of such systems are models with flat bands flat ; Flach . Flat-band in the one-magnon spectrum of the frustrated Heisenberg spin system means that magnons are localized in a small portion of the lattice (trapping cells). The localization of the magnons arises due to a destructive quantum interference, caused by the frustration. There are highly frustrated antiferromagnetic spin systems which supports a completely dispersionless magnon band flat ; shulen ; mak ; zhit , so that the excitations in this band are localized states. The localization of the one-magnon states is a base for the construction of multi-magnon states, because a state consisting of independent (non-overlapping) localized magnons is an exact eigenstate. Such systems include, for example, the delta-chain, the kagome lattice, kagome-like chains, the Tasaki lattice etc. An important feature of them is the triangular geometry of antiferromagnetic bonds. The ground state of such systems at the saturation magnetic field consists of independent localized magnons and it is macroscopically degenerated flat . The ground state and low-temperature properties for the antiferromagnetic Heisenberg models with flat band have been actively studied over last decades. It was shown that flat band physics may lead to new interesting phenomena such as the residual entropy at the saturation magnetic field, the zero-temperature magnetization plateau and the magnetization jump, an extra low-temperature peak in the specific heat etc. schmidt ; honecker ; shulen ; hon ; Derzhko . Such systems have enhanced magnetocaloric effect and are promising for efficient low-temperature cooling schmidt ; honecker ; zhitomir ; Derzhko .
@@
Recently it was found that the localized magnon states can exist in a certain frustrated spin system with competing ferro- (F) and antiferromagnetic (AF) interactions. An example of such model is spin-$\frac{1}{2}$ Heisenberg delta-chain with the ferromagnetic $J_{1}<0$ and the antiferromagnetic $J_{2}>0$ interactions (see Fig.1). For $J_{2}=-\frac{J_{1}}{2}$ the localized states exist in this model. However, in contrast with the AF-AF delta-chain, the localized magnon states of the F-AF chain are exact ground states at zero magnetic field. Besides, the ground state manifold contains the special states with overlapping magnons (localized multi-magnon complexes). So, the ground state degeneracy in this model is even higher than for AF-AF delta chain. The properties of such F-AF delta chain have been studied in Refs.KDNDR .
@@
In this paper we consider one more example of the model, in which the localized magnons lead to the macroscopic degeneracy of the ground state. This is the spin-$\frac{1}{2}$ Heisenberg diamond chain, shown in Fig.2. Generally, this model is one of examples where the interplay of quantum effects and frustration leads to a wide variety of the ground state phases and it has attracted a lot of attention in both experimental and theoretical studies. It was proposed as minimal theoretical model of several copper-based compounds as azurite $Cu_{3}(CO_{3})_{2}(OH)_{2}$ azur and alumoklyuchevskite $K_{3}Cu_{3}AlO_{2}(SO_{4})_{4}$ alum . Besides, the diamond model is also interesting in its own right. The ground state phase diagram of the diamond chain has been extensively studied in many works Takano ; Hon ; Sak ; Sakai ; Hu ; Krupn ; Gu ; Okam ; Strec ; Morita and it was shown that the diamond model exhibit various quantum phases depending on exchange interactions. In particular, the $1/3$ magnetization plateau was determined for certain ratios between exchange interactions Kitazawa ; Kab ; Okam ; Gu ; Morita .
@@
The paper is organized as follows. In Sec.II the model Hamiltonian with generally different exchange interactions is presented and the conditions for interactions in which the ground state has the macroscopic degeneracy are determined. It will be shown that, under certain condition, along with the localized one-magnon states there are also localized two- and three-magnon states. We present four types of the diamond chain with different ground state degeneracy. In the first type only one-magnon localized states exist. In the second one there are one-, two- and three- magnons in the trapping cell and this case includes the diamond chain with equivalent opposite sides of a diamond (we name it as the ‘distorted’ diamond chain). The third model is the ideal (symmetric) diamond chain and in the fourth type of the model one-magnon states are localized on diagonals of diamonds. The ground state degeneracy for these four types of diamond chain model are studied in Subsecs. IIA, IIB, IIC and IID. In concluding Section we give a summary of our results.
II Diamond chain in the critical point
@@
It is known KDNDR that the macroscopic ground state degeneracy in frustrated spin systems with competing (F) and (AF) interactions arises if the local Hamiltonian has several degenerate ground states, one of which is the state with the maximal spin. Four spins of diamond form one quintet ($S=2$), three triplets ($S=1$) and two singlets ($S=0$). The condition that the quintet and one of three triplets are degenerate with zero energy is
@@
Relations (3) and (4) determine the range of interaction values corresponding to the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases of model (1). For simplicity we chose the interaction $J_{2}$ as a parameter defining the transition point. Eq.(3) gives the critical value of this parameter. For $J_{2}$ less than the value given by Eq.(3), the ground state of model (1) is ferromagnetic. Further we propose that both relations (3) and (4) are satisfied.
@@
The energy of the ferromagnetic state of $\hat{H}$ with maximal total spin $S_{\max}=\frac{3n}{2}$ is zero. Therefore, the inequality in (5) turns in an equality and the ground state energy is zero.
The relations (3) and (4) can also be obtained alternatively from the condition that an one-magnon excitation band with the spin $S^{z}=S_{\max}^{z}-1$ is dispersionless (flat band) with zero energy. The dispersionless one-magnon states correspond to localized states which can be chosen as
@@
The states $\hat{\varphi}_{1i}\left|F\right\rangle$ are localized in the trapping cells between neighboring diamonds as shown in Fig.2 by shaded area. All $n$ states (6) are the ground states and linearly independent, because each $\hat{\varphi}_{1i}$ include the corresponding operator $s_{i}^{-}$, which is not shared with other functions $\hat{\varphi}_{1j}$.
@@
As follows from Eqs.(7),(8), there is one special case in which the denominator in these equations diverges:
@@
In this special case one-magnon states (6) are localized on diagonals of diamonds and can be written as
@@
Then, two- and three-magnon states (12) and (13) also belong to the ground state manifold. It is worth noting that when relation (14) is valid, the ground state degeneracy of each local Hamiltonian $\hat{H}_{i}$ increases. In this case, one singlet state is added to 8 degenerate ground states with $S=2$ and $S=1$.
@@
Relation (14) is satisfied automatically for important case of the ‘distorted’ diamond chain, in which $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but $J_{1}\neq J_{3}$).
@@
There is one more special case, when both relations (10) and (14) are satisfied simultaneously. In this case $J_{1}=J_{3}=J_{4}=$ $J_{5}=-1$ and $J_{2}=1$ (ideal diamond chain) and localized one-magnon functions are diagonal singlets $(\eta_{i}^{-}-\xi_{i}^{-})\left|F\right\rangle$. As will be shown below, the ground state degeneracy is maximal in this case.
@@
Figure 3: Examples of four different types of the diamond spin chain with macroscopic ground state degeneracy. The values of exchange integrals are written near the corresponding bonds. For clarity, the ferromagnetic bonds (all with $J=-1$) are shown by solid lines, antiferromagnetic bonds - by dashed lines.
@@
Thus, we have determined the necessary conditions (3) and (4) for the macroscopic ground state degeneracy of the diamond chain model (1). Based on the above analysis we can distinguish four different types of the diamond chain with the macroscopic ground state degeneracy, depending on whether the relations (10) and (14) are satisfied or not. Examples of four types of diamonds are shown in Fig.3. We will name these types as follows: the ‘general case’ is when none of the relations (10) and (14) are fulfilled (Fig.3a); the ‘distorted diamond case’ is when the relations (14) is satisfied, but Eq.(10) is not (Fig.3b); the ‘ideal diamond case’ is when both relations are fulfilled (Fig.3c); and the ‘diagonal diamond case’ is when the relations (10) is satisfied, but Eq.(14) is not (Fig.3d).
@@
Figure 4: Phase diagram of spin diamond chain (1) with diagonal bond interaction $J_{2}$ given by Eq.(3). The special cases are shown by solid line (‘distorted diamond case’), dashed line (‘diagonal state case’) and by solid circle (‘ideal diamond case’). Thick open circles denotes four particular cases of the diamond model shown in Fig.3 and signed with the corresponding letter a, b, c and d.
@@
As follows from the relations (10) and (14), it is convenient to plot the phase diagram of model (1) in the $J_{3}/J_{1}$ and $J_{5}/J_{4}$ axes (see Fig.4). In Fig.4 the ‘diagonal state case’ is shown by dashed line, the distorted diamond case by solid line and the ideal case by solid circle. The entire plane, except for these special cases, corresponds to the general case. However, in order to satisfy the conditions (4), there is a prescription for how to choose the signs of interactions $J_{i}$ corresponding to a definite point on the phase diagram Fig.4. If the point is in the region $J_{3}/J_{1}>0$, then the condition $J_{1}+J_{3}<0$ leads uniquely to the conclusion that both $J_{3}$ and $J_{1}$ are ferromagnetic ($J_{3}$, $J_{1}<0$). If $-1<J_{3}/J_{1}<0$, then $J_{3}>0$ and $J_{1}<0$. And, finally, if $J_{3}/J_{1}<-1$, then $J_{3}<0$ and $J_{1}>0$. Similar conditions are applied for the correct choice of signs of interactions $J_{4}$ and $J_{5}$ in different regions of the phase diagram. To illustrate, four particular cases of the diamond model shown in Fig.3, are represented by thick open circles in Fig.4 and signed with the corresponding letter: a, b, c and d.
@@ II.1 General case
Let us first consider the case when both relations (10) and (14) are not satisfied. The example of the diamond of this type is shown in Fig.3a, where the exchange integrals are chosen so that the diagonal interaction $J_{2}=0$. In the general case there are only one-magnon localized states with zero energy as in the F-AF delta-chain at the critical value of frustration parameter. In the latter the one-magnon state with zero energy is localized in a valley between neighboring triangles (see Fig.1). It is evident that the product of the independent (isolated) magnons belong to the ground state. As it is shown in KDNDR for the F-AF delta-chain, in addition to the independent localized states there are also specifically overlapping magnons in the ground states. For the general case of the diamond chain the problem of counting of these states is completely similar to that for the F-AF delta-chain, which was solved in KDNDR . Rewriting these results to the diamond chain with periodic boundary conditions (PBC) we obtain that the number of the ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is
@@
According to this equation the ground state degeneracy is exponentially large. It leads to the residual entropy $s_{0}=\frac{\ln W}{n}$ and $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$.
The properties of the excitation spectrum of the diamond chain in the general case are similar to that for the F-AF delta-chain. Numerical calculations show that the gap in the spectrum is exponentially small at $n\gg 1$ as in the F-AF delta-chain. The low-temperature thermodynamic properties in this case are determined by these low-lying excitations and they lead to specific behavior of thermodynamic quantities at low temperature. For example, $C(T)$ has low-temperature maximum. The magnetization curve calculated with the partition function accounting only the ground state manifold is
@@
As follows from this equation, the magnetization at $\frac{h}{T}\to 0$ tends to $\frac{1}{3}$. Therefore, the diamond chain in the general case is magnetically ordered in the transition point at $T=0$ and, therefore, the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
@@
Let us consider the ‘distorted’ diamond chain in which the trapping cell can contain localized one-, two- and three-magnon states. In this case condition (14) is satisfied but (10) is not. Further in this subsection we focus on the most interesting from the experimental point of view case $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but assuming $J_{1}\neq J_{3}$). The particular case of such choice corresponds to the distorted diamond chain shown in Fig.3b. In fact, this is a generic case for the the ‘distorted’ diamond chain, so that the results for the case $J_{3}\neq J_{4}$, $J_{1}\neq J_{5}$ are qualitatively the same.
It is again evident that the product of the independent (isolated) magnons are the ground states. However, similar to the previous case such states do not exhaust all possible ones with zero energy. For example, definite combinations of overlapping magnons are also the exact ground states. For the F-AF delta chain with open boundary condition (OBC) such multi-magnon states can be chosen as products
@@
As it was shown in KDNDR , the product (19) is the exact eigenfunction of each triangle Hamiltonian including $m$-th and $(m+1)$-th triangles.
For the diamond chain the analog of the operator $\Phi_{1}(m)$ has form
@@
For the diamond chain with two- and three-magnons localized in the same trapping cell this approach can be extended to include functions similar to $\Phi_{1}(m)$ for two- and three-magnon states. These functions have forms
@@
It can be checked that functions $\Phi_{\sigma}(m)\left|F\right\rangle$ ($\sigma=1,2,3$) are exact eigenfunctions of all local Hamiltonians with zero energy. Then we can construct for the diamond model with (OBC) the products
@@
These products are the eigenfunctions with zero energy of each local Hamiltonians and they are the ground states. We can interpret these products as the states of the system of $n$ non-interacting spins $\frac{3}{2}$. For example, let us consider the product
@@
This product can be identified as the state of the system with spin $s=\frac{3}{2}$ consisting of $(n-4)$ spins with $s^{z}=\frac{3}{2}$, one spin with $s^{z}=\frac{1}{2}$, two spins with $s^{z}=-\frac{1}{2}$ and one spin with $s^{z}=-\frac{3}{2}$. Generally, the number of the ground states of the diamond model in the spin sector $S^{z}$ coincides with that for the system of non-interacting spins $\frac{3}{2}$.
Further, we consider the case of the periodic boundary conditions (PBC). It is evident that a state consisting of isolated magnons
@@
and $S^{-}(i-1,i)$ is the total lowering operator of ($i-1)$-th and $i$-th diamonds. It could be check that the functions of the type (28) are the eigenstates with zero energy of the local Hamiltonians $\hat{H}_{i-1}$, $\hat{H}_{i}$, and all others.
We conjecture that similar to the OBC case the problem of counting number of the ground state is reduced to that for the system of $n$ non-interacting spins $\frac{3}{2}$. Then, the number of the ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is determined by equations
@@
where $(n-l),n_{1},n_{2},n_{3}$ are numbers of spins $s=\frac{3}{2}$ with $s_{z}=\frac{3}{2}$, $\frac{1}{2}$, $-\frac{1}{2}$, $-\frac{3}{2}$, correspondingly. The solution of a combinatorial problem (32) gives the number of the ground states $W(n,k)$ in form
@@
Results (33) and (34) have been confirmed by ED calculation of finite diamond chains. The residual entropy per spin $s$ is

### v2

@@ I Introduction
The low-dimensional quantum magnets on geometrically frustrated lattices have been extensively studied during the last years [1, 2]. There is a peculiar class of frustrated quantum magnets in which the ground state is macroscopically degenerate. Systems with macroscopic degeneracy have attracted great interest and are usually described by models with flat bands [3, 4]. Flat-band in the one-magnon spectrum of the frustrated Heisenberg spin system means that the magnons are localized in a small portion of the lattice (trapping cells). The localization of the magnons arises due to a destructive quantum interference, caused by frustration. There are highly frustrated antiferromagnetic spin systems which support a completely dispersionless magnon band [4, 5, 6, *zhit2, 8], so that the excitations in this band are localized states. The localization of one-magnon states is a base for the construction of multi-magnon states, because a state consisting of independent (non-overlapping) localized magnons is an exact eigenstate. Such systems include, for example, the delta-chain, the kagome lattice, kagome-like chains, the Tasaki lattice etc. Their important feature is the triangular geometry of antiferromagnetic bonds. The ground state of such systems at the saturation magnetic field consists of independent localized magnons and it is macroscopically degenerate [4]. The ground state and low-temperature properties for the antiferromagnetic Heisenberg models with flat band have been actively studied over the last decades. It was shown that flat band physics may lead to new interesting phenomena such as the residual entropy at the saturation magnetic field, the zero-temperature magnetization plateau and the magnetization jump, an extra low-temperature peak in the specific heat etc. [9, 10, 11, 8, 12]. Such systems have enhanced magnetocaloric effect and are promising for efficient low-temperature cooling [10, 11, 13, 12].
@@
Recently it was found that the localized magnon states can exist in a certain frustrated spin system with competing ferro- (F) and antiferromagnetic (AF) interactions. An example of such model is spin-$\frac{1}{2}$ Heisenberg delta-chain with the ferromagnetic $J_{1}<0$ and the antiferromagnetic $J_{2}>0$ interactions (see Fig.1). For $J_{2}=-\frac{J_{1}}{2}$ the localized states exist in this model. However, in contrast to the AF-AF delta-chain, the localized magnon states of the F-AF chain are exact ground states at zero magnetic field. Besides, the ground state manifold contains the special states with overlapping magnons (localized multi-magnon complexes). Thus, the ground state degeneracy in this model is even higher than the one for the AF-AF delta chain. The properties of such F-AF delta chain have been studied in Refs.[14, 15, *DKRS2, *DKRS3].
@@
In this paper we consider one more example of the model, in which the localized magnons lead to the macroscopic degeneracy of the ground state. This is the spin-$\frac{1}{2}$ Heisenberg diamond chain, shown in Fig.2. Generally, this model is one of the examples where the interplay of quantum effects and frustration leads to a wide variety of ground state phases; it has attracted a lot of attention in both experimental and theoretical studies. It was proposed as a minimal theoretical model of several copper-based compounds as azurite $Cu_{3}(CO_{3})_{2}(OH)_{2}$ [18] and alumoklyuchevskite $K_{3}Cu_{3}AlO_{2}(SO_{4})_{4}$ [19]. Besides, the diamond model is also interesting in its own right. The ground state phase diagram of the diamond chain has been extensively studied in many works [20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and it was shown that the diamond model exhibits various quantum phases depending on exchange interactions. In particular, the $1/3$ magnetization plateau was determined for certain ratios between exchange interactions [30, 31, 28, 27, 29].
@@
The paper is organized as follows. In Sec.II the model Hamiltonian with generally different exchange interactions is presented and the conditions for the interactions in which the ground state has macroscopic degeneracy are determined. It will be shown that, under certain conditions, along with the localized one-magnon states there are also localized two- and three-magnon states. We present four types of the diamond chain with different ground state degeneracy. In the first type only one-magnon localized states exist. In the second one there are one-, two- and three- magnons in the trapping cell and this case includes the diamond chain with equivalent opposite sides of a diamond (we name it as the ‘distorted’ diamond chain). The third model is the ideal (symmetric) diamond chain and in the fourth type of the model one-magnon states are localized on diagonals of diamonds. The ground state degeneracy for these four types of diamond chain models are studied in Subsecs. IIA, IIB, IIC and IID. In concluding Section we give a summary of our results.
II Diamond chain at the critical point
@@
It is known [14] that the macroscopic ground state degeneracy in frustrated spin systems with competing (F) and (AF) interactions arises if the local Hamiltonian has several degenerate ground states, one of which is the state with the maximal spin. Four spins of diamond form one quintet ($S=2$), three triplets ($S=1$) and two singlets ($S=0$). The condition that the quintet and one of three triplets are degenerate with zero energy is
@@
Relations (3) and (4) determine the range of interaction values corresponding to the transition point between the ferromagnetic and other (singlet or ferrimagnetic) ground state phases of model (1). For simplicity we chose the interaction $J_{2}$ as a parameter defining the transition point. Eq.(3) gives the critical value of this parameter. For $J_{2}$ less than the value given by Eq.(3), the ground state of model (1) is ferromagnetic. Further we assume that both relations (3) and (4) are satisfied.
@@
The energy of the ferromagnetic state of $\hat{H}$ with maximal total spin $S_{\max}=\frac{3n}{2}$ is zero. Therefore, the inequality in (5) turns into an equality and the ground state energy is zero.
The relations (3) and (4) can also be obtained alternatively from the condition that a one-magnon excitation band with the spin $S^{z}=S_{\max}^{z}-1$ is dispersionless (flat band) with zero energy. The dispersionless one-magnon states correspond to localized states which can be chosen as
@@
The states $\hat{\varphi}_{1i}\left|F\right\rangle$ are localized in the trapping cells between neighboring diamonds as shown in Fig.2 by a shaded area. All $n$ states (6) are the ground states and they are linearly independent, because each $\hat{\varphi}_{1i}$ includes the corresponding operator $s_{i}^{-}$, which is not shared with other functions $\hat{\varphi}_{1j}$.
@@
As follows from Eqs.(7),(8), there is one special case in which the denominator in these equations goes to zero:
@@
In this special case the one-magnon states (6) are localized on diagonals of diamonds and can be written as
@@
Then, two- and three-magnon states (12) and (13) also belong to the ground state manifold. It is worth noting that when the relation (14) is valid, the ground state degeneracy of each local Hamiltonian $\hat{H}_{i}$ increases. In this case, one singlet state is added to 8 degenerate ground states with $S=2$ and $S=1$.
@@
Relation (14) is satisfied automatically for the important case of the ‘distorted’ diamond chain, in which $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but $J_{1}\neq J_{3}$).
@@
There is one more special case, when both relations (10) and (14) are satisfied simultaneously. In this case $J_{1}=J_{3}=J_{4}=$ $J_{5}=-1$ and $J_{2}=1$ (ideal diamond chain) and localized one-magnon functions are diagonal singlets $(\eta_{i}^{-}-\xi_{i}^{-})\left|F\right\rangle$. As it will be shown below, the ground state degeneracy is maximal in this case.
@@
Figure 3: Examples of four different types of diamond spin chains with macroscopic ground state degeneracy. The values of the exchange integrals are written close to the corresponding bonds. For clarity, the ferromagnetic bonds (all with $J=-1$) are shown by solid lines, antiferromagnetic bonds - by dashed lines.
@@
Thus, we have determined the necessary conditions (3) and (4) for the macroscopic ground state degeneracy of the diamond chain model (1). Based on the above analysis we can distinguish four different types of the diamond chain with the macroscopic ground state degeneracy, depending on whether the relations (10) and (14) are satisfied or not. Examples of four types of diamonds are shown in Fig.3. We will name these types as follows: the ‘general case’ is the one when none of the relations (10) and (14) are fulfilled (Fig.3a); the ‘distorted diamond case’ is realized when the relation (14) is satisfied, but Eq.(10) is not (Fig.3b); the ‘ideal diamond case’ is when both relations are fulfilled (Fig.3c); and the ‘diagonal diamond case’ is when the relation (10) is satisfied, but Eq.(14) is not (Fig.3d).
@@
Figure 4: Phase diagram of spin diamond chain (1) with diagonal bond interaction $J_{2}$ given by Eq.(3). The special cases are shown by solid line (‘distorted diamond case’), dashed line (‘diagonal state case’) and by solid circle (‘ideal diamond case’). Thick open circles denote four particular cases of the diamond model shown in Fig.3 and they are signed with the corresponding letters a, b, c and d.
@@
As it follows from the relations (10) and (14), it is convenient to plot the phase diagram of the model (1) in the $J_{3}/J_{1}$ and $J_{5}/J_{4}$ axes (see Fig.4). In Fig.4 the ‘diagonal state case’ is shown by a dashed line, the distorted diamond case by a solid line and the ideal case by a solid circle. The entire plane, except for these special cases, corresponds to the general case. However, in order to satisfy the conditions (4), there is a prescription on how to choose the signs of interactions $J_{i}$ corresponding to a definite point on the phase diagram Fig.4. If the point is in the region $J_{3}/J_{1}>0$, then the condition $J_{1}+J_{3}<0$ leads uniquely to the conclusion that both $J_{3}$ and $J_{1}$ are ferromagnetic ($J_{3}$, $J_{1}<0$). If $-1<J_{3}/J_{1}<0$, then $J_{3}>0$ and $J_{1}<0$. And, finally, if $J_{3}/J_{1}<-1$, then $J_{3}<0$ and $J_{1}>0$. Similar conditions are applied for the correct choice of signs of interactions $J_{4}$ and $J_{5}$ in different regions of the phase diagram. To illustrate, four particular cases of the diamond model shown in Fig.3, are represented by thick open circles in Fig.4 and signed with the corresponding letters: a, b, c and d.
@@ II.1 General case
Let us first consider the case when both relations (10) and (14) are not satisfied. The example of the diamond of this type is shown in Fig.3a, where the exchange integrals are chosen so that the diagonal interaction $J_{2}$ is zero. In the general case there are only one-magnon localized states with zero energy as in the F-AF delta-chain at the critical value of the frustration parameter. In the latter case the one-magnon state with zero energy is localized in a valley between neighboring triangles (see Fig.1). Any state consisting of independent (isolated) magnons has zero energy and belongs to the ground state manifold. As it is shown in [14] for the F-AF delta-chain, in addition to the independent localized states there are also specifically overlapping magnons in the ground states. For the general case of the diamond chain the problem of counting of these states is completely similar to the one for the F-AF delta-chain, which was solved in [14]. Rewriting these results to the case of the diamond chain with periodic boundary conditions (PBC) we find that the number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is
@@
According to this equation the ground state degeneracy is exponentially large. It leads to the finite residual entropy per spin (total number of spins is $N=3n$)
$s_{0}=\frac{\ln W(n)}{N}$ (18)
and $s_{0}=\frac{1}{3}\ln 2$ at $n\gg 1$.
Figure 5: Dependence of the energy gap (logarithmic scale) on the system size for the general case (Fig.3a), calculated numerically for $N=12,18,24$ (exact diagonalization and Lanczos algorithm).
The properties of the excitation spectrum of the diamond chain in the general case are similar to that for the F-AF delta-chain. Numerical calculations show that the gap in the spectrum is exponentially small at $n\gg 1$ as in the F-AF delta-chain (see Fig.5).
Figure 6: Temperature dependencies of the specific heat (solid line), entropy (dashed line) and the product of magnetic susceptibility and temperature, $\chi T$, (dotted line) for the general case (Fig.3a), calculated numerically for $N=18$ (exact diagonalization). For better visibility the specific heat is multiplied by a factor of 3, $3C(T)$.
The low-temperature thermodynamic properties in this case are determined by these low-lying excitations and they lead to the specific behavior of the thermodynamic quantities at low temperature. For example, $C(T)$ has low-temperature maximum and a long low-temperature tail, as shown in Fig.6. The entropy in Fig.6 tends to the finite value in accordance with Eq.(18) at $T\to 0$ and increases to the value $s_{0}=\ln 2$ at $T\to\infty$.
The behavior of the magnetization and the uniform susceptibility of the diamond chain (in the general case) is similar to the behavior of the F-AF delta-chain, and this point was discussed in detail in [14]. We give here a brief summary of the results [14].
Since there are exponentially low excited states, the contribution of excited states to magnetization and susceptibility can not be neglected even at very low temperatures. Nevertheless, it is useful to consider the magnetization given only by the contribution of the ground state manifold. This magnetization $m_{0}$ is a function of variable $x=\frac{h}{T}$. At $x\gg\frac{1}{N}$ the magnetization $m_{0}$ for the general case of diamond chain is
@@
Figure 7: Magnetization curve as a function of $h/T$ for the general case (Fig.3a), calculated numerically for $N=18$ and $T=0.001$ (thick solid line) and given by Eq.(19) (thin solid line).
The magnetization curve (19) is shown in Fig.7 together with the magnetization curve calculated numerically for $N=18$ and $T=0.001$. As one can see in Fig.7, the calculated curve coincides with the curve (19) for $h/T>1$, but the limit $h/T\to 0$ is different. The reason of this discrepancy is related to the singular behavior of magnetization and susceptibility in low magnetic fields.
The numerically calculated magnetization is zero at $h=0$ and increases linearly with increasing magnetic field, $m=\chi h$. The magnetic susceptibility $\chi$ depends on the temperature and the system size. An analysis similar to the one given in [14] shows that in the thermodynamic limit the magnetic susceptibility diverges at $T\to 0$ as $\chi\sim T^{-\alpha}$ with the exponent $\alpha\approx 1.25$. Thus, the true magnetization curve in the thermodynamic limit, $m(h)$, starts at $m(0)=0$, then increases linearly with a very steep slope,
$m\sim\frac{h}{T^{\alpha}}$ (20)
so that it approaches the value $m=1/3$ at $h_{c}\sim T^{\alpha}$. For $h>h_{c}$ the true magnetization is described by Eq.(19).
When the magnetization curve is plotted as a function of $h/T$, as shown in Fig.7, the low-field region where the true magnetization curve deviates from Eq.(19) is limited by $h_{c}/T\sim T^{\alpha-1}$, so that this region vanishes in the limit $T\to 0$. This implies that the magnetization at $\frac{h}{T}\to 0$ tends to $\frac{1}{3}$. Therefore, the diamond chain in the general case is magnetically ordered at the transition point at $T=0$ and, therefore, the magnetization undergoes a jump from $m=\frac{1}{2}$ in the ferromagnetic phase to $m=\frac{1}{3}$ at the transition point.
@@
Let us consider the ‘distorted’ diamond chain in which the trapping cell can contain localized one-, two- and three-magnon states. In this case condition (14) is satisfied but (10) is not. Further in this subsection we focus on the most interesting from the experimental point of view case $J_{3}=J_{4}$ and $J_{1}=J_{5}$ (but assuming $J_{1}\neq J_{3}$). A particular case of such choice corresponds to the distorted diamond chain shown in Fig.3b. In fact, this is a generic case for the the ‘distorted’ diamond chain, so that the results for the case $J_{3}\neq J_{4}$, $J_{1}\neq J_{5}$ are qualitatively similar.
All possible states consisting of independent (isolated) magnons have zero energy and belong to the ground states. However, in similarity to the previous case such states do not exhaust all possible ones with zero energy. For example, definite combinations of overlapping magnons are also the exact ground states. For the F-AF delta chain with open boundary condition (OBC) such multi-magnon states can be chosen as products
@@
As it was shown in [14], the product (21) is the exact eigenfunction of each triangle Hamiltonian including $m$-th and $(m+1)$-th triangles.
For the diamond chain the analog of the operator $\Phi_{1}(m)$ has the form
@@
For the diamond chain with two- and three-magnons localized in the same trapping cell this approach can be extended to include functions similar to $\Phi_{1}(m)$ for two- and three-magnon states. These functions have the forms
@@
It can be checked that the functions $\Phi_{\sigma}(m)\left|F\right\rangle$ ($\sigma=1,2,3$) are exact eigenfunctions of all local Hamiltonians with zero energy. Then we can construct the following products for the diamond model with (OBC)
@@
These products are the zero-energy eigenfunctions of each local Hamiltonian and they form the ground states. We can interpret these products as the states of the system of $n$ non-interacting spins $\frac{3}{2}$. For example, let us consider the product
@@
This product can be identified as the state of the system with spin $s=\frac{3}{2}$ consisting of $(n-4)$ spins with $s^{z}=\frac{3}{2}$, one spin with $s^{z}=\frac{1}{2}$, two spins with $s^{z}=-\frac{1}{2}$ and one spin with $s^{z}=-\frac{3}{2}$. Generally, the number of ground states of the diamond model in the spin sector $S^{z}$ coincides with that for the system of non-interacting spins $\frac{3}{2}$.
Further, we consider the case of the periodic boundary conditions (PBC). Any state consisting of isolated magnons
@@
and $S^{-}(i-1,i)$ is the total lowering operator of ($i-1)$-th and $i$-th diamonds. It can be checked that the functions of the type (30) are the eigenstates with zero energy of the local Hamiltonians $\hat{H}_{i-1}$, $\hat{H}_{i}$, and of all the others.
We conjecture that in similarity to the OBC case the problem of counting the number of ground states is reduced to the one for a system of $n$ non-interacting spins $\frac{3}{2}$. Then, the number of ground states $W(n,k)$ in the spin sector $S^{z}=S_{\max}-k$ is determined by equations
@@
where $(n-l),n_{1},n_{2},n_{3}$ are numbers of spins $s=\frac{3}{2}$ with $s_{z}=\frac{3}{2}$, $\frac{1}{2}$, $-\frac{1}{2}$, $-\frac{3}{2}$, correspondingly. The solution of a combinatorial problem (34) gives the number of ground states $W(n,k)$ in form
@@
Results (35) and (36) have been confirmed by ED calculation of finite diamond chains. The residual entropy per spin in accordance with Eq.(18) is